40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Apr 10 12:17:34 2019, Anjali, Update, 2micronLasers, Intensity noise stabilization of 2 micron source 2_micron_setup.jpgLaser_chara_setup.pngLaser_characteristics.pdfAOM_chara_setup.pngAOM_characteristics.pdf
    Reply  Wed Apr 10 18:58:48 2019, rana, Update, 2micronLasers, Intensity noise stabilization of 2 micron source 
       Reply  Wed Apr 24 09:01:49 2019, Anjali, Update, 2micronLasers, Intensity noise stabilization of 2 micron source Intensity_stabilisation.png
Message ID: 2323     Entry time: Wed Apr 10 12:17:34 2019     Reply to this: 2325
Author: Anjali 
Type: Update 
Category: 2micronLasers 
Subject: Intensity noise stabilization of 2 micron source 

[Aidan, Anjali]

  • Attachment #1 shows the cleaned up setup for  2 micron experiment
  • We also have received the AOM . The initial test is to do the intensity stabilisation of 2 micron laser diode using AOM
  • Today, we started with the laser and AOM characterisation
  • Attachment #2 shows the schematic of the experimental setup for the laser characterisation. Laser output is connected to a Faray isolator and the isolator output is measured using Thorlabs thermal power meter.
  • The current limit on the current controller is set to 95 mA (maximum allowed current to the laser diode is 120 mA) and thermistor is set to 10 kohm (maximum value of thermistor resistance is 10 kohm)
  • Attachment #3 shows the output power as a function of input current to the laser diode.
  • From the data sheet of the laser diode, the threshold current is 20 mA.For a given inut current,  the output power measured is less compared to the value given in the data sheet. This is beacuse of the insertion loss of the isolator in the path.
  • The slope efficieny is calculted to be 0.018 mW/mA. The value given in the data sheet is 0.3 mW/mA. The reduction is because of the insertion loss of the isolator
  • The output of isolator is then connected to the AOM. Attachement #4 shows the schematic of the setup. The zeroth order port of the AOM is connected to the power meter and the output power from the first order port is blocked.
  • We then applied a dc voltage to the modulation input port of the AOM driver. We expect a decrease in the power from the zeroth order port with increase in the input voltage to the  AOM driver as the light is getting deflcted and power is geting coupled to the first order port.
  • Attachmet #5 shows the variation in power from zeroth order port with input voltage to the AOM driver.As expected, the power levels are decreasing with increase in input voltage
  • It is also observed that the output power levels are fluctuating. This could be because of temperature fluctuation and the thermal power meter is sensitive to that. 
Attachment 1: 2_micron_setup.jpg  1.280 MB  Uploaded Wed Apr 10 13:21:31 2019  | Hide | Hide all
2_micron_setup.jpg
Attachment 2: Laser_chara_setup.png  20 kB  Uploaded Wed Apr 10 16:07:47 2019  | Hide | Hide all
Laser_chara_setup.png
Attachment 3: Laser_characteristics.pdf  3 kB  Uploaded Wed Apr 10 16:08:00 2019  | Hide | Hide all
Laser_characteristics.pdf
Attachment 4: AOM_chara_setup.png  35 kB  Uploaded Wed Apr 10 16:08:20 2019  | Hide | Hide all
AOM_chara_setup.png
Attachment 5: AOM_characteristics.pdf  4 kB  Uploaded Wed Apr 10 16:08:32 2019  | Hide | Hide all
AOM_characteristics.pdf
ELOG V3.1.3-